OpenStudy (anonymous):

let (x',y') be the solution of the following equation: (2x)^ln2=(3y)^ln3 3^lnx=2^lny. then x' is ?

OpenStudy (anonymous):

2x^ln2 = 3y^ln3 which means that 3y = (2x)^(ln2-ln3) which means that 3y = (2x)^(ln(2/3)) or y = (2x)^(ln(2/3)) substitute in the other equation and solve for x

OpenStudy (anonymous):

sorry, y = ((2x)^ln(2/3))/3

OpenStudy (dumbcow):

i disagree, 2x^(ln2/ln3) does not equal 2x^(ln2-ln3) ->y = ((2x)^(ln2/ln3))/3

OpenStudy (dumbcow):

i get x' = 1/2, y' = 1/3

OpenStudy (anonymous):

oh, dumbcow is right, I made a mistake there